# HLSFactory: A Framework Empowering High-Level Synthesis Datasets for Machine Learning and Beyond

### v2

@@ V-A Case Study 1: Design Space Coverage
We then illustrate the qualitative coverage of the design space in Fig. 6. This space is the 2-D embedding space of HLS-reported and post-implementation metrics generated using a PacMAP [23] dimensional reduction. [^1] PaCMAP is one of many dimensional reduction techniques which can be used to visualize data in this manner. However, many dimensional reduction techniques have trade-offs and free parameters with no meaningful defaults. Each of the base designs is depicted as large emphasized points within this embedding space; sampled designs from the same base design (top panel) or the same benchmark (bottom panel) have matching colors. The convex hulls around same-colored points show the portion of the embedding space covered by design space expansion from each base design or benchmark. This clearly shows that sampling from the expanded design space results in non-overlapping coverage that otherwise would not appear in the final dataset.
Figure 6: Embedding of sampled designs across selected benchmarks. Base designs without optimizations are emphasized. Design points and locations are the same between both panels; they are only colored and grouped differently.
V-B Case Study 2: Speedup of Fine-Grained Design Parallelism
We evaluate our fine-grained parallelism strategy described in Sec. IV-C using a case study synthesizing (Vitis HLS) designs sampled from Polybench, MachSuite, and CHStone across 32 cores.
@@ V-B Case Study 2: Speedup of Fine-Grained Design Parallelism
Figure 7: Parallel execution of Vitis HLS synthesis. Top panel shows core utilization over time with naive parallelism across datasets; bottom panel shows our fine-grained design parallelism across datasets.
V-C Case Study 3: Targeting Different Vendors
@@
As described in Sec. III-B2, HLSFactory includes an OptDSL parser that recognizes Vitis HLS optimization directives in opt_template.tcl, such as the set_directive_unroll and set_directive_array_partition commands. We can therefore build our Intel-lowering frontend on top of this functionality.
@@
Figure 8: Distribution of post-implementation metrics for PolyBench and MachSuite designs ($n=1340$) using Intel’s HLS flow.
In total, our end-to-end Intel flow extends the HLSFactory user APIs in Table II with three Intel equivalents: OptDSLFrontendIntel(Frontend) as described, IntelHLSSynthFlow(ToolFlow) to invoke i++ for HLS, and IntelQuartusImplFlow(ToolFlow) to invoke Quartus for implementation. We run this flow on designs sampled from PolyBench and MachSuite and plot the resulting metrics in Fig. 8. Intel’s HLS tool does not report overall latency estimates, but it optimizes each kernel’s throughput by maximizing clock speed, which we use as a proxy for performance.
V-D Case Study 4: Adding Auxiliary Design Collections
Third-party researchers may have existing, synthesizable, vendor-specific HLS designs to integrate into HLSFactory, but they may not want or need to create an OptDSL specification for them.
For instance, the authors of LightningSim [24] collect 33 synthesizable open-source designs for AMD/Xilinx Vitis HLS to evaluate their simulation tool, including designs from AMD/Xilinx sample code repositories [25, 26], algorithm implementations from Kastner et al.’s Parallel Programming for FPGAs [27], and graph neural network implementations from FlowGNN [28]. These designs are all provided in a standard format, each having a Tcl script setup.tcl to set up a Vitis HLS project for synthesis.
@@ V-D Case Study 4: Adding Auxiliary Design Collections
Many other works [29, 30, 31, 32] were also easily integrated with HLSFactory in a similar fashion; the code is available online.
V-E Case Study 5: Integrating Released Data from Other Work
Figure 9: Distribution of HLS-estimated metrics from selected HLSFactory benchmarks (PolyBench, CHStone, and MachSuite; $n=167$) vs. HLSyn ($n=3371$).
@@ V-E Case Study 5: Integrating Released Data from Other Work
We illustrate how HLSFactory can integrate pre-generated data from prior works—in this case, HLSyn [16]. HLSyn provides both the source code (with places to template optimization directives) of their selected kernels, as well as associated metrics for HLS-reported resource usage and latency for sampled designs. We write a DataAggregator subclass to integrate this data into HLSFactory.
@@
The HLSyn flow is built on top of AutoDSE [33, 34] and the Merlin compiler [35, 36], both of which are open-source software tools aimed at extending high-level synthesis capabilities for optimized design space exploration (DSE) and source-to-source translations of HLS designs. These tools suggests future work to integrate AutoDSE and the Merlin compiler as custom flows in HLSFactory, allowing designs to be built from the design space specifications defined in AutoDSE and synthesized with the Merlin compiler.
V-F Case Study 6: Regression Benchmarking of HLS Tools
@@
The results are shown in Fig. 10. We show distributions of the tool runtime, HLS-estimated latency, LUT usage, and FF usage across tool versions. We also report the $p$-value for a paired two-tailed Wilcoxon signed-rank test [37, p. 350] [^2] We use the paired Wilcoxon signed-rank test because we cannot assume our metrics are normally distributed. The use of this test is often misrepresented in statistical education with discussion outside the scope of this work [38, 39, 40]. This test compares the “pseudo-medians” of two distributions; we also do not adjust for differences in their variances. We include statistical testing mainly to show how HLSFactory can be used for more in-depth statistical analysis alongside hardware design benchmarking. and indicate cases with a $p$-value less than $\alpha=0.05$ with an asterisk, indicating a statistically significant difference. Note that for certain metrics, the mean and median shift in opposite directions between tool versions.
Figure 10: Distribution of HLS tool metrics from two versions of Vitis HLS.
VI Conclusion
HLSFactory brings a much-needed principled approach to generating datasets of HLS designs. Our case studies show a small sample of what can be done when we have a flexible, reproducible way to generate data from HLS designs; we hope that our brief explorations show that there is substantial untapped potential for future research into how ML can be applied to HLS.

### v3

@@ 5.2. Case Study 2: Design Space Coverage
We then illustrate the qualitative coverage of the design space in Fig. 7. This space is the 2-D embedding space of HLS-reported and post-implementation metrics generated using a PacMAP (Wang et al., 2021) dimensional reduction. Each of the base designs is depicted as large emphasized points within this embedding space; sampled designs from the same base design (top panel) or the same benchmark (bottom panel) have matching colors. The convex hulls around same-colored points show the portion of the embedding space covered by design space expansion from each base design or benchmark. This clearly shows that sampling from the expanded design space results in non-overlapping coverage that otherwise would not appear in the final dataset.
Figure 6. Effect of design sampling to cover more design space. Sampled designs cover a wider range of metrics than base designs with no optimizations. Latency is HLS estimated; resources are post-implementation. Note that these are stacked density plots to show the effect of cumulative design sampling.
Figure 7. Embedding of sampled designs across selected benchmarks. Base designs without optimizations are emphasized. Design points and locations are the same between both panels; they are only colored and grouped differently.
Figure 8. Parallel execution of Vitis HLS synthesis. Top panel shows core utilization over time with naive parallelism across datasets; bottom panel shows our fine-grained design parallelism across datasets.
5.3. Case Study 3: Speedup of Fine-Grained Design Parallelism
We evaluate our fine-grained parallelism strategy described in Sec. 4.3 using a case study synthesizing designs sampled from Polybench, MachSuite, and CHStone using Vitis HLS across 32 CPU cores.
@@ 5.3. Case Study 3: Speedup of Fine-Grained Design Parallelism
5.4. Case Study 4: Targeting Different Vendors
@@
As described in Sec. 3.2.2, HLSFactory includes an OptDSL parser that recognizes Vitis HLS optimization directives in opt_template.tcl, such as the set_directive_unroll and set_directive_array_partition commands. We can therefore build our Intel-lowering frontend on top of this functionality.
@@
Figure 9. Distribution of post-implementation metrics for PolyBench and MachSuite designs ($n=1340$) using Intel’s HLS flow.
In total, our end-to-end Intel flow extends the HLSFactory user APIs in Table 2 with three Intel equivalents: OptDSLFrontendIntel(Frontend) as described above, IntelHLSSynthFlow(ToolFlow) to invoke i++ for HLS, and IntelQuartusImplFlow(ToolFlow) to invoke Quartus for implementation. We run this flow on designs sampled from PolyBench and MachSuite and plot the resulting metrics in Fig. 9. Intel’s HLS tool does not report overall latency estimates, but it optimizes each kernel’s throughput by maximizing clock speed, which we use as a proxy for performance.
5.5. Case Study 5: Adding Auxiliary Design Collections
Third-party researchers may have existing, synthesizable, vendor-specific HLS designs to integrate into HLSFactory, but they may not want or need to create an OptDSL specification for them. For instance, the authors of LightningSim (Sarkar and Hao, 2023) collect 33 synthesizable open-source designs for AMD/Xilinx Vitis HLS to evaluate their simulation tool, including designs from AMD/Xilinx sample code repositories (AMD/Xilinx, 2021, 2022), algorithm implementations from Kastner et al.’s Parallel Programming for FPGAs (Kastner et al., 2018), and graph neural network implementations from FlowGNN (Sarkar et al., 2023a). These designs are all provided in a standard format, each having a Tcl script setup.tcl to set up a Vitis HLS project for synthesis.
@@ 5.5. Case Study 5: Adding Auxiliary Design Collections
Many other works (Chen and Hao, 2023, 2022; Zhang et al., 2020; Sarkar et al., 2023b) were also easily integrated with HLSFactory in a similar fashion; the code is available online.
5.6. Case Study 6: Integrating Released Data from Other Works
Figure 10. Distribution of HLS-estimated metrics from selected HLSFactory benchmarks (PolyBench, CHStone, and MachSuite; $n=167$) vs. HLSyn ($n=3371$).
@@ 5.6. Case Study 6: Integrating Released Data from Other Works
We illustrate how HLSFactory can integrate pre-generated data from prior works—in this case, HLSyn (Bai et al., 2023). HLSyn provides both the source code (with places to template optimization directives) of their selected kernels, as well as associated metrics for HLS-reported resource usage and latency for sampled designs. We write a DataAggregator subclass to integrate this data into HLSFactory.
@@
The HLSyn flow is built on top of AutoDSE (Sohrabizadeh et al., [n. d.]; aut, [n. d.]) and the Merlin compiler (Cong et al., [n. d.]; mer, [n. d.]), both of which are open-source software tools aimed at optimized design space exploration (DSE) and source-to-source translation. These tools suggest future work to integrate AutoDSE and the Merlin compiler as custom flows in HLSFactory, allowing designs to be built from the design space specifications defined in AutoDSE and synthesized with the Merlin compiler.
Figure 11. Distribution of HLS tool metrics from two versions of Vitis HLS.
5.7. Case Study 7: Regression Benchmarking HLS Synthesis Tools
@@
The results are shown in Fig. 11. We show distributions of the tool runtime, HLS-estimated latency, LUT usage, and FF usage across tool versions. We also report the $p$-value for a paired two-tailed Wilcoxon signed-rank test (Conover, 1999, p. 350) and indicate cases with a $p$-value less than $\alpha=0.05$ with an asterisk, indicating a statistically significant difference. Note that for certain metrics, the mean and median shift in opposite directions between tool versions.
6. Conclusion
HLSFactory brings a much-needed principled approach to generating datasets of HLS designs. Our case studies show a small sample of what can be done when a flexible, reproducible way to generate data from HLS designs is available. We demonstrate that there is substantial untapped potential for future research into how ML can be applied to HLS.
@@ 6. Conclusion
Acknowledgements
This research was supported in part by National Science Foundation (NSF) Grant #2326894, NVIDIA Applied Research Accelerator Program Grant, and the Texas Advanced Computing Center (TACC). Any opinions, findings, conclusions, or recommendations are those of the authors and not of the funding agencies. We also thank Georgia Tech Research Institute for direct funding of selected authors.